Question 1: Large country trade:
Country A: demand: Q=400-P, supply: Q=P-20 Country B: demand: Q=300-P, supply: Q=2P-30

  1. Which country is importing? What is the global price under free trade? (10%)

  2. Compute the social surplus of each country. (5%)

  3. If the importing country impose a $20 tariff, what is the change in social surplus in

    each country? (15%)
Answer #1

Country A equilibrium:

400-P=P-20

2P=420

P=210

Q=190

Country B equilibrium:

300-P=2P-30

3P=330

P=110

Q=190

Since, price in country A is higher. Country A is importing.

-------------------------------------------------------------------------------------------------------------------------------------------------

Under free trade: global demand = Q = 400-P + 300-P = 700-2P

global supply = Q=P-20+2P-30 = 3P-50

Global equilibrium: 700-2P = 3P-50

5P=750

P=150

Q=400

Thus, global price under free trade = 150

----------------------------------------------------------------------------------------------------------------------------------------------------------

At P=150, domestic demand in country A = Q = 400-P = 400-150 = 250

But domestic demand is 190, thus, social surplus = 250-190 = 60

At P=150, domestic demand in country B = Q=300-P = 300-150 = 150

But domestic demand is 190, thus social surplus = 150-190 =-40

------------------------------------------------------------------------------------------------------------------------------------------------------------

Now tariff = $20

Global demand = Q = 400-P+20+300-P = 720-2P

Global Supply =  Q=P-20+2P-30 = 3P-50

Global equilibrium: 720-2P=3P-50

5P = 770

P=154

Q=372

At P=154, domestic demand in country A = Q = 400-P = 400-154 = 246

But domestic demand is 190, thus, social surplus = 246-190 = 56

At P=154, domestic demand in country B = Q=300-P = 300-154 = 146

But domestic demand is 190, thus social surplus = 146-190 =-44